### v1

@@ 2.2 Optimal model, barycentre model, and average drift model
$={\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}\left[\frac{{\rm d}{\mathbb{Q}}^{\mathfrak{b}}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\,\frac{e^{-\eta_{0}\int_{0}^{T}g_{t}\,{\rm d}t-\eta_{1}\,f_{T}}}{{\mathbb{E}}^{{}^{{\mathbb{Q}}^{\mathfrak{b}}}}\left[\,e^{-\eta_{0}\int_{0}^{T}g_{t}\,{\rm d}t-\eta_{1}\,f_{T}}\right]}\frac{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}{{\rm d}{\mathbb{Q}}^{\mathfrak{b}}}\int_{0}^{T}g_{u}\,{\rm d}u\right]$
@@
Similarly, we can rewrite the second constraint expectation as
@@
By 2.4, when $\eta=\eta^{*}$ the Equations (14) and (15) vanish and hence, under the measure ${\mathbb{Q}}[\theta_{\eta^{*}}]$ the constraints are binding. Replacing $\eta$ with $\eta^{*}$ in (12) yields the representation of the RN derivative $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ in the statement.
We can view the optimal measure as a distortion of the barycentre measure by incorporating the constraints. Specifically, alternatively to optimisation problem (P), we can in a first step find the barycentre and then in a second step distort the barycentre model to incorporate the constraints. This idea is formalised in the next statement. Modifying stochastic processes to include expectation and running cost constraints has been studied in [15] for Lévy-Itô processes, and in [18] for pure jump processes in a financial risk management setting.
@@
Let Assumptions 2.1, 2.2, 2.3, and 2.4 be satisfied. Then, optimisation problem (P) is equivalent to
@@
${\mathcal{Q}}^{\mathfrak{b}}:=\Bigg{\{}\;{\mathbb{Q}}[\theta]~{}\Big{|}~{}$ $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}^{{}^{\mathfrak{b}}}}=\exp\big{\{}-\tfrac{1}{2}\int_{0}^{T}|\lambda(t,X_{t})|^{2}{\rm d}t-\int_{0}^{T}\lambda(t,X_{t})\,{\rm d}W_{t}^{{\mathfrak{b}}}\big{\}}\,,$
$\text{where}\quad\lambda(t,x):=\sigma^{-1}(t,x)\big{(}\mu^{{\mathfrak{b}}}(t,x)-\theta(t,x)\big{)}\,,$
$\text{and}\quad{\mathbb{E}}^{{\mathbb{Q}}^{{\mathfrak{b}}}}\left[\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}^{{}^{\mathfrak{b}}}}\right]=1\ \Bigg{\}}\,,$
@@
where $\eta^{*}$ is the solution to Equation (10). Recall that the solution to optimisation problem (P), ${\mathbb{Q}}[\theta_{\eta^{*}}]$, is given by Proposition 2.9. Next, using (13) we obtain
@@
Moreover, all the results in this section, in particular Proposition 2.2, 2.4, Proposition 2.9, and Proposition 2.11, hold when replacing $\eta_{0}g(\cdot,\cdot)$ with $\boldsymbol{\eta_{0}}\cdot\mathfrak{g}(\cdot,\cdot)$, where $\boldsymbol{\eta_{0}}:=(\eta_{0}^{1},\ldots\eta_{0}^{I})$ and $\mathfrak{g}:=(g^{1},\ldots,g^{I})$, and similarly when replacing $\eta_{1}f(\cdot)$ with $\boldsymbol{\eta_{1}}\cdot\mathfrak{f}(\cdot)$, where $\boldsymbol{\eta_{1}}:=(\eta_{1}^{1},\ldots\eta_{1}^{J})$ and $\mathfrak{f}:=(f^{1},\ldots,f^{J})$, and where $\cdot$ denotes the dot product.
@@ 3 Deep learning algorithms
In low dimensions, such as $d=1$ or $2$, we can approximate the solution to optimisation problem (P) using finite-difference methods by (a) solving for the optimal Lagrange multipliers by solving the appropriate PDE, and (b) once the optimal Lagrange multipliers are obtained, using a second finite-difference scheme to approximate $\omega(t,x)$ which then provides, through Proposition 2.2, the optimal drift. In higher dimensions, however, finite-difference and other PDE methods are not applicable. Hence, to overcome the curse of dimensionality, we propose two deep learning algorithms to approximate the solution to optimisation problem (P). The first deep learning algorithm approximates the drift of a candidate measure and uses the difference in the RN densities as a loss function. The second deep learning algorithm approximates a transformation of the value function, specifically $\omega(t,x)=e^{-L_{\eta^{*}}(t,x)}$, by leveraging the notion of elicitability.
@@ 3.1 Learning the optimal drift
In this section, we develop a deep learning approach that directly learns the stochastic exponential that drives the measure change from ${\mathbb{Q}}[{\bar{\mu}}]$ to the optimal measure ${\mathbb{Q}}[\theta_{\eta^{*}}]$. For this, we write the measure change from a candidate measure ${\mathbb{Q}}[\theta]$ to ${\mathbb{Q}}[{\bar{\mu}}]$ as
(16) $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\exp\left\{-\tfrac{1}{2}\int_{0}^{T}|{\bar{\lambda}}(t,X_{t})|^{2}\,{\rm d}t-\int_{0}^{T}{\bar{\lambda}}(t,X_{t})\,{\rm d}{\overline{W}}_{t}\right\},$
where ${\bar{\lambda}}(t,x):=\sigma^{-1}(t,x)({\bar{\mu}}(t,x)-\theta(t,x))$ and $({\overline{W}}_{t})_{{t\in[0,T]}}$ are ${\mathbb{Q}}[{\bar{\mu}}]$-Brownian motions. To approximate the optimal measure, we parameterise the drift process (under the candidate measure ${\mathbb{Q}}[\theta]$) $(\theta_{t})_{{t\in[0,T]}}$ by a neural network with parameters ${\mathfrak{a}}$, and write this as $\theta[{\mathfrak{a}}](t,X_{t})$ – which takes $t$ and $X_{t}$ as inputs and outputs the drift at time $t$ and state $X_{t}$.
The approach we take to learning the drift is as follows. First, introduce a time grid $\{t_{0},\dots,t_{N}\}$, where $t_{i}:=i\Delta t$ and $\Delta t:=\frac{T}{N}$. Second, simulate Brownian motions under the measure ${\mathbb{Q}}[{\bar{\mu}}]$, denoted as ${\overline{W}}_{t_{0}},{\overline{W}}_{t_{1}},\dots,{\overline{W}}_{t_{N}}$ and write $\Delta{\overline{W}}_{t_{i}}:={\overline{W}}_{t_{i+1}}-{\overline{W}}_{t_{i}}$. Third, use an Euler–Maruyama discretisation of $X$ via
$X_{t_{i}}=X_{t_{i-1}}+{\bar{\mu}}_{t_{i-1}}\,\Delta t+\sigma_{t_{i-1}}\,\Delta{\overline{W}}_{t_{i}}\,,$
where ${\bar{\mu}}_{t_{i-1}}:={\bar{\mu}}(t_{i-1},X_{t_{i-1}})$ and $\sigma_{t_{i-1}}:=\sigma(t_{i-1},X_{t_{i-1}})$. From these sample paths, for the current estimate of the parameters ${\mathfrak{a}}$, we then obtain samples of $\theta_{t_{i}}^{\mathfrak{a}}:=\theta[{\mathfrak{a}}](t_{i},X_{t_{i}})$ and evaluate
@@
Note that the above quantities can be approximated from samples under the fixed measure ${\mathbb{Q}}[{\bar{\mu}}]$. The algorithm for finding $\eta^{*}$ is presented in Algorithm 1. Note that while 2.4 gives $\eta^{*}$ as a solution to a set of non-linear equations, the expectations are with respect to the barycentre measure ${\mathbb{Q}}^{\mathfrak{b}}$. Thus, to utilise 2.4, we need samples paths under ${\mathbb{Q}}^{\mathfrak{b}}$, which require to solve for the barycentre model. The above approach makes use of the readily available average drift model, which is straightforwardly obtained from simulations of the experts’ models.
@@
${\mathfrak{L}}\big{[}\theta[{\mathfrak{a}}]\big{]}:={\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\log\left(\frac{{\rm d}{\mathbb{Q}}\big{[}\theta[{\mathfrak{a}}]\big{]}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)-\log\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)\right]\,,$
with $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ given in Proposition 2.9. All terms in the loss function may be approximated by simulating under the single measure ${\mathbb{Q}}[{\bar{\mu}}]$. We update the parameters ${\mathfrak{a}}$ using gradient descent, via ${\mathfrak{a}}\leftarrow{\mathfrak{a}}-r_{\mathfrak{a}}\,\nabla_{\mathfrak{a}}{\mathfrak{L}}\big{[}\theta[{\mathfrak{a}}]\big{]}$, where $r_{\mathfrak{a}}$ is a learning rate. Further details are given in Algorithm 2.
@@
2generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}},\ldots,{\overline{W}}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(X_{t_{0}}^{{[m]}},\ldots,X_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
@@
${\hat{f}}[\eta]:=\frac{1}{M}\sum_{m\in{\mathfrak{M}}}f\big{(}X_{t_{N}}^{{[m]}}\big{)}\,\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)^{{[m]}}\quad\text{and}\quad{\hat{g}}[\eta]:=\frac{1}{M}\sum_{\begin{subarray}{c}m\in{\mathfrak{M}}\\ i=0,\ldots,N-1\end{subarray}}g\big{(}t_{i},X_{t_{i}}^{{[m]}}\big{)}\,\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)^{{[m]}}\,\Delta t;$
@@
4       generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}}\dots\overline{W}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(X_{t_{0}}^{{[m]}}\dots X_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
@@
6      approx. $\int_{0}^{T}|{\bar{\lambda}}_{u}|^{2}{\rm d}u$ and $\int_{0}^{T}{\bar{\lambda}}_{u}^{\intercal}{\rm d}{\overline{W}}$ and use (16) to generate samples $\Big{(}\Big{(}\frac{{\rm d}{\mathbb{Q}}\big{[}\theta[{\mathfrak{a}}]\big{]}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\big{)}^{{[m]}}\Big{)}_{m\in{\mathfrak{M}}}$;

### v2

@@ 4.1 Optimal model
Similarly, we can rewrite the second constraint as
@@
By 4.1, when $\eta=\eta^{*}$ the Equations (75) and (76) vanish and hence, under the measure ${\mathbb{Q}}[\theta_{\eta^{*}}]$ the constraints are binding. Replacing $\eta$ with $\eta^{*}$ in (73) yields the representation of the RN derivative $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ in the statement of the proposition.
4.2 Distorting the barycentre model
We can view the optimal measure as a distortion of the barycentre measure by incorporating the constraints. Specifically, alternatively to optimisation problem (P), we can in a first step find the barycentre and then in a second step distort the barycentre model to incorporate the constraints. Figure 2 provides a visualisation of this idea. The left panel shows that we may obtain the barycentre model and then impose the constraints from the single barycentre model or directly find the optimal model by searching within the set of constraints and minimizing the weighted KL divergence.
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{1}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{2}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{3}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{4}}}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}^{\mathfrak{b}}}$$\mathcal{C}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}^{*}}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{1}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{2}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{3}}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{{\mathbb{P}}_{4}}}$$\mathcal{C}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{\mathbb{Q}}^{*}}$
Figure 2: Illustration that first finding the barycenter model and imposing constraints is equivalent to directly imposing the constraints and minimizing the weighted KL divergence.
This idea is formalised in the next statement. Modifying stochastic processes to include expectation and running cost constraints has been studied in [18] for Lévy–Itô processes, and in [20] for pure jump processes in a financial risk management setting.
@@ 4.2 Distorting the barycentre model
Let the assumptions in Theorem 4.3 and 4.1 hold. Then, optimisation problem (P) is equivalent to
@@
${\mathcal{Q}}^{\mathfrak{b}}:=\Bigg{\{}\;{\mathbb{Q}}[\theta]~{}\Big{|}~{}$ $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}^{{\mathfrak{b}}}}=\exp\left\{-\tfrac{1}{2}\int_{0}^{T}|\lambda_{t}^{\mathfrak{b}}|^{2}{\rm d}t-\int_{0}^{T}\lambda_{t}^{{\mathfrak{b}}\intercal}\,{\rm d}W_{t}^{{\mathfrak{b}}}\right\}\,,$
$\text{where}\quad\lambda_{t}^{\mathfrak{b}}:=\sigma^{-1}_{t}\big{(}\mu^{{\mathfrak{b}}}_{t}-\theta_{t}\big{)}\,,$
[…]
$\qquad\text{and}\quad{\mathbb{E}}^{{\mathbb{Q}}^{{\mathfrak{b}}}}\left[\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}^{{}^{\mathfrak{b}}}}\right]=1\ \Bigg{\}}\,,$
@@
where $\eta^{*}$ is the solution to Equation (65). Recall that the solution to optimisation problem (P), ${\mathbb{Q}}[\theta_{\eta^{*}}]$, is given by Proposition 4.7.
Next, using (74) we obtain
@@
Moreover, all the results in this section, in particular Theorem 4.3, Proposition 4.7, and Proposition 4.9, hold when replacing $\eta_{0}g(\cdot,\cdot)$ with $\boldsymbol{\eta_{0}}\cdot\mathfrak{g}(\cdot,\cdot)$, where $\boldsymbol{\eta_{0}}:=(\eta_{0}^{1},\ldots\eta_{0}^{I})$ and $\mathfrak{g}:=(g^{1},\ldots,g^{I})$, and similarly when replacing $\eta_{1}f(\cdot)$ with $\boldsymbol{\eta_{1}}\cdot\mathfrak{f}(\cdot)$, where $\boldsymbol{\eta_{1}}:=(\eta_{1}^{1},\ldots\eta_{1}^{J})$ and $\mathfrak{f}:=(f^{1},\ldots,f^{J})$, and where $\cdot$ denotes the dot product.
@@ 5 Deep learning algorithms
In low dimensions, such as $d=1$ or $2$, we can approximate the solution to optimisation problem (P) using finite-difference methods by (a) solving for the optimal Lagrange multipliers by solving the appropriate PDE, and (b) once the optimal Lagrange multipliers are obtained, using a second finite-difference scheme to approximate $\omega(t,x)$ which then provides, through (70), the optimal drift. In higher dimensions, however, finite-difference and other traditional PDE methods run into difficulties due to the curse of dimensionality. Hence, to overcome the curse of dimensionality, we propose two deep learning algorithms to approximate the solution to optimisation problem (P). The first deep learning algorithm approximates the drift of a candidate measure and uses the difference in the RN densities as a loss function. The second deep learning algorithm approximates a transformation of the value function, specifically $\omega(t,x)=e^{-L_{\eta^{*}}(t,x)}$, by leveraging the notion of elicitability — which allows one to estimate conditional statistics of a distribution function, e.g., a conditional expectation, via a strictly convex optimization problem. Often when elicitability is used to estimate conditional statistics, a parametric assumption on the dependence of the conditioning variables are made, e.g., the dependence is assumed to be linear, quadratic, or polynomial in the conditioning variables. Rather, when utilizing neural networks, due to their universal approximation ability, parametric assumptions are not necessary. While neural networks and deep learning approaches have been utilised in a variety in financial modelling contexts, the specific algorithms presented here are tailored to the constrained KL barycentre problem that we pose. In this sense, the algorithms are novel. We make use of the fact that neural networks are universal approximators, but tune the algorithms to providing approximations to our specific problem formulation.
@@ 5.1 Learning the optimal drift
In this section, we develop a deep learning approach that directly learns the stochastic exponential — the RN derivative — that drives the measure change from ${\mathbb{Q}}[{\bar{\mu}}]$ to the optimal measure ${\mathbb{Q}}[\theta_{\eta^{*}}]$. For this, we write the measure change from a candidate measure ${\mathbb{Q}}[\theta]$ to ${\mathbb{Q}}[{\bar{\mu}}]$ as
(77) $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\exp\left\{-\tfrac{1}{2}\int_{0}^{T}|{\bar{\lambda}}(t,X_{t})|^{2}\,{\rm d}t-\int_{0}^{T}{\bar{\lambda}}(t,X_{t})^{\intercal}\,{\rm d}{\overline{W}}_{t}\right\},$
where ${\bar{\lambda}}(t,x):=\sigma^{-1}(t,x)({\bar{\mu}}(t,x)-\theta(t,x))$ and $({\overline{W}}_{t})_{{t\in[0,T]}}$ are ${\mathbb{Q}}[{\bar{\mu}}]$-Brownian motions. To approximate the optimal measure, we parametrise the drift process (under the candidate measure ${\mathbb{Q}}[\theta]$) $(\theta_{t})_{{t\in[0,T]}}$ by a neural network with parameters ${\mathfrak{a}}$, and write this as $\theta[{\mathfrak{a}}](t,X_{t})$ — which takes $t$ and $X_{t}$ as inputs and outputs the drift at time $t$ and state $X_{t}$.
The approach we take to learning the drift is as follows. First, introduce a time grid $\{t_{0},\dots,t_{N}\}$, where $t_{i}:=i\Delta t$, $i=0,\ldots,N$, and $\Delta t:=\frac{T}{N}$. Second, simulate sample paths of Brownian motions under the measure ${\mathbb{Q}}[{\bar{\mu}}]$, denoted as ${\overline{W}}_{t_{0}},{\overline{W}}_{t_{1}},\dots,{\overline{W}}_{t_{N}}$ and write their increments as $\Delta{\overline{W}}_{t_{i}}:={\overline{W}}_{t_{i+1}}-{\overline{W}}_{t_{i}}$. Third, use an Euler–Maruyama discretisation of $X$, which we denote ${\tilde{X}}$, via
$\tilde{X}_{t_{i}}=\tilde{X}_{t_{i-1}}+{\bar{\mu}}_{t_{i-1}}\,\Delta t+\sigma_{t_{i-1}}\,\Delta{\overline{W}}_{t_{i}}\,,$
where ${\bar{\mu}}_{t_{i-1}}:={\bar{\mu}}(t_{i-1},\tilde{X}_{t_{i-1}})$ and $\sigma_{t_{i-1}}:=\sigma(t_{i-1},\tilde{X}_{t_{i-1}})$ — with a slight abuse of notation we do not place tildes on ${\bar{\mu}}$ and $\sigma$, despite them being evaluated along the Euler–Maruyama discretisation of $X$. From these sample paths, for the current estimate of the parameters ${\mathfrak{a}}$, we then obtain sample paths of $\theta[{\mathfrak{a}}]$ by evaluating the neural network for $\theta[{\mathfrak{a}}]$ along the sample paths of $\tilde{X}$ that we already obtained, i.e., by evaluating $\theta_{t_{i}}^{\mathfrak{a}}:=\theta[{\mathfrak{a}}](t_{i},\tilde{X}_{t_{i}})$. We then evaluate
@@
Note that the above quantities can be approximated from samples under the fixed measure ${\mathbb{Q}}[{\bar{\mu}}]$. The algorithm for finding $\eta^{*}$ is presented in Algorithm 1. Note that while 4.1 gives $\eta^{*}$ as a solution to a set of non-linear equations, the expectations are with respect to the barycentre measure ${\mathbb{Q}}^{\mathfrak{b}}$. Thus, to utilise 4.1, we need samples paths under ${\mathbb{Q}}^{\mathfrak{b}}$, which require to solve for the barycentre model. The above approach makes use of the readily available average drift model, which is straightforwardly obtained from simulations of the experts’ models.
@@
${\mathfrak{L}}\big{[}\theta[{\mathfrak{a}}]\big{]}:={\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\left(\log\left(\frac{{\rm d}{\mathbb{Q}}\big{[}\theta[{\mathfrak{a}}]\big{]}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)-\log\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)\right)^{2}\right]\,,$
with $\frac{{\rm d}{\mathbb{Q}}[\theta[{\mathfrak{a}}]]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ given in (77) and $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ given in Proposition 4.7. All terms in the loss function may be approximated by simulating under the single measure ${\mathbb{Q}}[{\bar{\mu}}]$. We update the parameters ${\mathfrak{a}}$ using gradient descent, via ${\mathfrak{a}}\leftarrow{\mathfrak{a}}-r_{\mathfrak{a}}\,\nabla_{\mathfrak{a}}{\mathfrak{L}}\big{[}\theta[{\mathfrak{a}}]\big{]}$, where $r_{\mathfrak{a}}$ is a learning rate. Further details are given in Algorithm 2.
@@
2generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}},\ldots,{\overline{W}}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(\tilde{X}_{t_{0}}^{{[m]}},\ldots,\tilde{X}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
@@
${\hat{f}}[\eta]:=\frac{1}{M}\sum_{m\in{\mathfrak{M}}}f\big{(}\tilde{X}_{t_{N}}^{{[m]}}\big{)}\,\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)^{{[m]}}\quad\text{and}\quad{\hat{g}}[\eta]:=\frac{1}{M}\sum_{\begin{subarray}{c}m\in{\mathfrak{M}}\\ i=0,\ldots,N-1\end{subarray}}g\big{(}t_{i},\tilde{X}_{t_{i}}^{{[m]}}\big{)}\,\left(\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\right)^{{[m]}}\,\Delta t;$
@@
4       generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}}\dots\overline{W}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(\tilde{X}_{t_{0}}^{{[m]}}\dots\tilde{X}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
@@
6      approx. $\int_{0}^{T}|{\bar{\lambda}}_{u}|^{2}{\rm d}u$ and $\int_{0}^{T}{\bar{\lambda}}_{u}^{\intercal}{\rm d}{\overline{W}}$ and use (77) to generate samples $\Big{(}\Big{(}\frac{{\rm d}{\mathbb{Q}}[\theta[{\mathfrak{a}}]]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}\big{)}^{{[m]}}\Big{)}_{m\in{\mathfrak{M}}}$;